Your company wants to raise $8.5 million by issuing 20-year zero-coupon bonds. If the yield to maturity on the bonds will be 9% (annual compounded APR), what total face value amount of bonds must you issue? The total face value amount of bonds that you must issue is s (Round to the nearest cent.)
Answer #1

The Face Value amount of the bond is ‘F’, then

8,500,000 = F/(1+9%)^20 (since it is a zero coupon bond)

F = 8,500,000 * (1+9%)^20 = $47,637,491.53